# TALE: A Tool-Augmented Framework for Reference-Free Evaluation of Large Language Models

### v1

@@ 5.2 Effect of iterations
We conducted an ablation study to evaluate the effect of the number of iterations in the TALE framework. Figure 2 shows that increasing the number of iterations generally improves the judge performance across candidate models on both AmbigQA and HotpotQA. Iteration 3, which serves as our default configuration, consistently provides the best trade-off between performance and computational cost. However, performance declines at iteration 4 across all models. This drop is likely caused by an overabundance of sources, resulting in the accumulation of redundant or irrelevant information, increased context length, and potential model confusion. These factors can dilute decision quality and lead to marginal or negative returns.
@@ 5.3 Effect of removing the query making
Table 4: Comparison of Macro F1 scores between Judge w/o Query (GPT-4o as-as-judge) and TALE (GPT-4o as-as-judge). Note that directly using the input question without generating refined queries is still considered a form of querying, but for clarity, we refer to this setting as w/o Query.
@@ 6 Conclusion
We presented TALE, a novel framework for evaluating LLMs. TALE integrates external evidence via an iterative process. Through extensive experiments, we found that TALE achieves substantial to perfect agreement with human annotators. TALE is also aligned with reference-based metrics such as F1. However, studying TALE compared to human evaluations, we found that TALE is a reliable alternative to reference-based metrics with interpretable evidence-aware evaluations. Regardless, TALE has limitations. Its short-term memory provides a persistent record of the evaluation process. While this enhances interpretability and traceability, it remains constrained by the model’s ability to process information within its context window during subsequent reasoning. Future work could explore integrating a long-term memory with a recall-based mechanism that selectively retrieves relevant traces from past evaluations, such as episodic and semantic memory (Park et al., 2023). In addition to this, TALE may inherit biases from external data sources, including sycophancy bias (Sharma et al., 2023), where the model may align with false claims instead of critically assessing them. Addressing these biases and ensuring robust evidence selection will be essential for broader deployment.
Acknowledgments
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), RGPIN-2022-03943, Canada Foundation of Innovation (CFI), and Research Nova Scotia. Advanced computing resources are provided by ACENET, the regional partner in Atlantic Canada, and the Digital Research Alliance of Canada.
@@ A.2 Prompting
We employ a template-based prompting strategy for both response generation and evaluation. For candidate models, we utilize few-shot Chain-of-Thought (CoT) prompts, incorporating 6 examples per dataset to encourage detailed, reasoning-driven, and structured responses (see Figure 3).
@@
Figure 3: Examples of few-shot Chain-of-Thought (CoT) prompts for candidate answer generation.
TALE prompts candidate LLMs using few-shot Chain-of-Thought (CoT) reasoning to elicit faithful and interpretable outputs across each module. Each module in the evaluation pipeline is guided by a carefully constructed prompt, often in a one-shot format, that combines role-playing instructions with explicit reasoning goals. Below, we describe the prompting strategy for each component.
@@ B.2 TALE with a small open-source model
B.5 TALE failure cases
@@
In some cases, TALE encounters conflicting evidence across multiple search iterations. While the framework is designed to iteratively refine its understanding and gather reliable information, it sometimes struggles to appropriately resolve contradictory information. Table 13 presents a scenario where the model failed to handle conflicting evidence when assessing the number of band members in Lush and P.O.D.
@@
Explanation Despite P.O.D.’s confirmed four-member structure, the model ignored evidence indicating Lush’s changing member count. It failed to account for the time-specific membership changes, which contradicted the claim that both bands "consist" of four members. This decision highlights TALE’s difficulty in resolving contradictory information.
Table 13: Example of conflicting evidence where TALE failed to resolve contradictions in band member counts.

### v2

@@ 5.2 Effect of iterations
We conducted an ablation study to evaluate the effect of the number of iterations in TALE. Figure 2 shows that increasing the number of iterations generally improves the judge performance across candidate models on both AmbigQA and HotpotQA. Iteration 3, which serves as our default configuration, consistently provides the best trade-off between performance and computational cost. However, performance declines at iteration 4 across all models. This drop is likely caused by an overabundance of sources, resulting in the accumulation of redundant or irrelevant information, increased context length, and potential model confusion. These factors can dilute decision quality and lead to marginal or negative returns.
@@ 5.3 Effect of removing the query making
Table 4: Macro F1 scores between Judge w/o Query and TALE. Note that directly using the input question without generating refined queries is still considered a form of querying, but for clarity, we refer to this setting as w/o Query.
5.4 Cost analysis
With three iterations, TALE issues three Serper queries and consumes 4,410 input plus 989 output LLM tokens per answer. Using GPT-4o-mini pricing (\$0.15 per $10^{6}$ input tokens and \$0.60 per $10^{6}$ output tokens) together with Serper pricing (\$0.30 per $10^{3}$ queries), this amounts to \$0.00126 (LLM) + \$0.00090 (search) = \$0.00216 per item. Evaluating all 300 AmbigQA instances therefore costs \$0.65 and, at 27 s per instance, completes in about 2.3 h. Three human annotators need 15 h (300 items $\times$ 3 min) and \$225 (at \$15 h-1). Thus TALE is ~6.7$\times$ faster and ~350$\times$ cheaper while retaining human-level accuracy ($\kappa\approx 0.91$).
@@ 6 Conclusion
We presented TALE, a framework for evaluating LLMs. TALE integrates external evidence via an iterative process. Through experiments, we found that TALE achieves substantial to perfect agreement with human annotators. TALE is also aligned with reference-based metrics such as F1. However, studying TALE compared to human evaluations, we found that TALE is a reliable alternative to reference-based metrics with interpretable evidence-aware evaluations. Regardless, TALE has limitations. Its short-term memory provides a persistent record of the evaluation process. While this enhances interpretability and traceability, it remains constrained by the model’s ability to process information within its context window during subsequent reasoning. Future work could explore integrating a long-term memory with a recall-based mechanism that selectively retrieves relevant traces from past evaluations, such as episodic and semantic memory (Park et al., 2023). In addition to this, TALE may inherit biases from external data sources, including sycophancy bias (Sharma et al., 2023), where the model may align with false claims instead of critically assessing them. Addressing these biases and ensuring robust evidence selection will be essential for broader deployment.
@@ A.2 Prompting
We employ a template-based prompting strategy for both response generation and evaluation. For candidate models, we utilize few-shot Chain-of-Thought (CoT) prompts (Gou et al., 2024), incorporating 6 examples per dataset to encourage detailed, reasoning-driven, and structured responses (see Figure 3).
@@
Figure 3: Examples of few-shot CoT (Gou et al., 2024) prompts for candidate answer generation.
TALE prompts candidate LLMs using few-shot CoT reasoning to elicit faithful and interpretable outputs across each module. Each module in the evaluation pipeline is guided by a carefully constructed prompt, often in a one-shot format, that combines role-playing instructions with explicit reasoning goals. Below, we describe the prompting strategy for each component.
@@ B.2 TALE with a small open-source model
Although Mistral 7B exhibits limitations in complex reasoning and instruction following, it still achieves meaningful alignment with human judgments and surpasses no-tool baselines. Since evaluation typically runs offline, users can select a judge model that balances performance and resource constraints. Furthermore, TALE’s model-agnostic design ensures that it remains compatible with ongoing improvements in smaller open-source LLMs, enhancing its accessibility and long-term utility.
@@
Appendix C Additional ablations and analysis
This section presents three extensions to our study. First, we compare TALE with three stronger non-iterative baselines to evaluate its relative performance. Second, we assess the framework’s robustness to prompt variations. Third, we analyze failure cases to better understand the limitations and potential areas for improvement in TALE.
C.1 Comparison with stronger non-iterative baselines
We evaluate three non-iterative baselines that exclude TALE’s iterative summarise–reflect–refine process but retain the judgment prompt. These methods help disentangle the contributions of tool augmentation, sampling-based self-consistency, and model diversity.
Across all baselines, we use GPT-4o to generate the candidate answer. For the Single-Pass Tool-Augmented Judge and Self-Consistency settings, GPT-4o also serves as the judge model. In the Multi-LLM Majority Voting setup, GPT-4o is the candidate model, while three different judge models—GPT-4o, GPT-3.5, and Mistral 7B—independently evaluate the same input to support diverse model reasoning.
C.1.1 Single-Pass tool-augmented judge
To isolate the effect of tool access alone, we introduce a baseline that issues exactly one web query and performs no further reasoning or iteration. Given a question $q$ and candidate answer $\hat{y}$, the judge LLM formulates a single Serper search, retrieves the top-3 snippets, and immediately produces a true/false verdict with rationale. All prompt instructions are identical to those in the full TALE pipeline; the only change is the removal of the summarise–reflect–refine loop. This baseline, therefore, quantifies the benefit of a one-shot evidence grab, sitting midway between the Judge without Tool-Augmentation and the full iterative framework.
As shown in Table 9, on AmbigQA, this naive tool baseline raises Cohen’s $\kappa$ from 0.381 (judge without tools) to 0.658, confirming that one round of external evidence already yields a substantial improvement. However, TALE’s pushes $\kappa$ further to 0.914. A similar pattern appears on HotpotQA, where $\kappa$ climbs from 0.375 (no tool) to 0.583 (single-pass) and then to 0.701 with TALE.
C.1.2 Self-consistency judge
For Self-Consistency (Wang et al., 2023), we let the GPT-4o judge sample 10 independent verdicts at temperature 0.7 and take a simple majority. In self-consistency, we excluded an external tool or reference answer, so the model relies exclusively on its parametric knowledge.
Table 9 shows that self-consistency gives a modest boost over the vanilla judge (no tool), but it still lags far behind the tool-augmented baselines, confirming that external evidence, especially when gathered iteratively is critical for reliable objective judgment.
C.1.3 Multi-LLM majority voting
Inspired by PoLL (Verga et al., 2024) and DAFE (Badshah & Sajjad, 2025), we instructed three different LLMs—GPT-4o, GPT-3.5, and Mistral 7B—with the same question and candidate answer, asking each to provide a True/False verdict along with a brief explanation. By applying majority voting across the three model outputs, we determined the final answer. This ensemble approach leverages the diverse reasoning patterns of different models, increasing overall robustness and reducing individual model biases.
This approach produces $\kappa$ = 0.609 on AmbigQA and 0.527 on HotpotQA, outperforming self-consistency but remaining below the single-pass tool ($\kappa$ = 0.658/0.583) and far below TALE ($\kappa$ = 0.914/0.701).
Task No Tool Self-Consistency Multi-LLMs Single-Pass TALE
AmbigQA 0.381 0.515 0.609 0.658 0.914
HotpotQA 0.375 0.493 0.527 0.583 0.701
Table 9: Cohen’s $\kappa$ agreement between LLM-based evaluators and the human majority vote. TALE achieves the highest alignment with human judgments across both AmbigQA and HotpotQA. Single-Pass denotes tool use without iteration and TALE components (i.e., summarize, reflect, and refine).
C.2 Prompt variations
To evaluate TALE’s robustness to prompt design, we re-ran the full framework (with GPT-4o as both candidate and judge) under three prompt variants:
  • Original (1-shot CoT): Our default configuration, which includes detailed instructions and a single CoT example per module.
  • Few-shot CoT: An extended version of the prompt containing three CoT examples instead of one.
  • Simplified (0-shot w/o CoT): Prompt that removes CoT examples for every module.
As shown in Table 10, the few-shot prompt yields a modest improvement over the original 1-shot setup. Removing CoT examples entirely results in a performance drop of at most 3 percentage points in both Cohen’s $\kappa$ and macro-F1. Interestingly, the default configuration remains highly stable: re-running TALE with the original prompt reproduced the same $\kappa$ and F1 scores as reported in prior experiments, underscoring its consistency.
Task Prompt Variant Cohen’s $\kappa$ Macro-F1
AmbigQA Original (1-shot CoT) 0.914 0.957
Few-shot CoT 0.920 0.960
Simplified (0-shot) 0.882 0.932
HotpotQA Original (1-shot CoT) 0.701 0.850
Few-shot CoT 0.706 0.853
Simplified (0-shot) 0.672 0.824
Table 10: TALE performance under prompt variations. The default (1-shot CoT) is robust, while removing CoT examples yields a small but consistent performance drop.
C.3 TALE failure cases
@@
At each round, TALE’s reflector module takes the aggregated evidence summary and generates a chain-of-thought reflection, pointing out which attributes seem well-supported, which appear contradictory or missing, and where further detail is needed (Section 3.3).
The next query is generated from that reflection, so contradictions are an explicit signal to search for clarifying evidence. We do not hard-code a credibility score; instead, TALE relies on cross-source agreement and iterative follow-up. All retrieved snippets and their domains will be released so that future work can plug in credibility weighting without altering the loop.
@@
At each round, TALE’s reflector module takes the aggregated evidence summary and generates a CoT reflection, pointing out which attributes seem well-supported, which appear contradictory or missing, and where further detail is needed. The next query is generated from that reflection, so contradictions are an explicit signal to search for clarifying evidence. However, in some cases, TALE encounters conflicting evidence across multiple search iterations. It sometimes struggles to appropriately resolve contradictory information. Table 15 presents a scenario where the model failed to handle conflicting evidence when assessing the number of band members in Lush and P.O.D.
@@
Table 15: Example of conflicting evidence where TALE failed to resolve contradictions in band member counts. Despite P.O.D.’s confirmed four-member structure, the model ignored evidence indicating Lush’s changing member count. It failed to account for the time-specific membership changes, which contradicted the claim that both bands "consist" of four members. This decision highlights TALE’s difficulty in resolving contradictory information.
Topic drift in query refinement.
To analyze how often TALE’s query refinement step introduces topic drift or irrelevant queries, we examined this within the 100 instances from our error analysis. Since our TALE configuration uses up to 3 iterations (1 initial query, up to 2 refined queries), this involved inspecting the query generation process across these stages.
We found that (see Table 16):
  • Initial queries, generated directly from the input question, were generally on-topic (sufficient) and relevant to the core question being asked. However, due to the ambiguity of some questions, we noticed some off-topic queries in this stage.
  • Topic drift, while not the most dominant error, did occur during the refinement stages (iterations 2 and 3). Out of the 200 refined queries analyzed, only 6 queries in the second iteration and 7 in the third iteration (a total of 13 out of 200, or 6.5%) were judged “insufficient” (indicative of topic drift or clear irrelevance).
Query Category Initial Query 2nd Query 3rd Query
Sufficient (On-topic) 77 69 70
Partially Sufficient 15 25 22
Insufficient (Off-topic) 8 6 7
Table 16: Topic drift across 300 queries (100 instances, 3 queries per instance = 300).
Appendix D Limitations and future work
Although TALE significantly improves the reliability and accuracy of LLM evaluation in dynamic scenarios, it has several limitations.
Dependency on judge LLM capabilities.
TALE’s absolute performance is influenced by the strength of the underlying judge model. While our experiments indicate graceful degradation with smaller models like Mistral 7B, the best performance requires stronger models such as GPT-4o.
Diminishing returns with iterative refinement.
Our empirical analysis shows that performance gains plateau and sometimes degrade beyond a certain number of iterative evidence-gathering steps due to redundant information and increased strain on the LLM context window. Although we currently select an optimal iteration count empirically, future research can develop adaptive stopping criteria, leveraging real-time evidence sufficiency metrics to halt iterations dynamically and efficiently.
Context window constraints.
While our short-term memory strategy reduces token input, limitations inherent in current LLM context windows still pose challenges, especially for longer evaluation sequences. Future implementations should investigate episodic memory that selectively offloads less relevant information, retrieving only critical snippets required for context-aware judgment, thus more effectively managing context length.
Source bias and quality control.
Although TALE’s reflector module actively identifies inconsistencies across multiple evidence sources, our current implementation does not explicitly quantify source credibility. To mitigate the impact of source biases and misinformation, future work should prioritize the integration of credibility scoring or weighting mechanisms that dynamically assess evidence reliability.
Addressing these limitations will enhance TALE’s robustness, scalability, and effectiveness, making it even more valuable for evaluating LLM outputs in diverse, real-world scenarios.
Appendix E Ethical considerations
TALE aims to enhance evaluation transparency and factual accuracy by grounding judgments in externally retrieved, verifiable evidence. However, several ethical considerations require attention.
Reliance on external sources.
The evidence quality in TALE depends significantly on the retrieved external content, which may introduce unintended biases or inaccuracies. Although our iterative refinement explicitly mitigates biases by cross-checking multiple sources, TALE does not currently include explicit credibility scoring. Researchers and practitioners should remain cautious about inherent biases in external sources and proactively manage credibility assessments.
Human oversight and accountability. TALE substantially improves automated LLM evaluation accuracy, but human oversight remains critical, especially in high-stakes decision-making contexts. It is imperative to recognize the inherent limitations of automated evaluations, maintain human accountability and critical review, particularly when evaluating sensitive or consequential content.
Computational costs and accessibility. TALE’s iterative, tool-augmented process significantly increases computational costs and inference latency compared to simpler evaluation methods, as explicitly shown in our analysis. Although this computational overhead is justified by gains in accuracy for offline evaluation scenarios such as benchmarking and auditing, researchers must consider the environmental and resource accessibility implications. We encourage future research to optimize computational efficiency and ensure broad accessibility.
By continuously addressing these ethical dimensions, TALE can support more reliable, transparent, and socially responsible AI evaluation practices.